OpenStudy (skittles_for_life6422):

AWARD WILL BE GIVEN AND WILL FAN !!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!METAL!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! a ball is thrown into the air with an initial upward velocity of 60 ft/s. its height (h) in feet after t seconds is given by the function h=-16t^2+60t+6. what will the height be at t=3 seconds?

OpenStudy (skittles_for_life6422):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

To find the heigh of the ball at t=3 seconds, we sumply plug 3 into our h(t)= -16t^2+60t+6 h(3)= -16*(3^2)+60*3+6 h(3)= -16*9+180+6 h(3)= 42 At t=3 seconds, the ball is 42 feet above the ground.

OpenStudy (alexandervonhumboldt2):

@skittles_for_life6422

OpenStudy (skittles_for_life6422):

THANK YOU A MILLION TIMES HERE IS YOU METAL I ALREADY FANNED YOU

OpenStudy (skittles_for_life6422):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

thx a lot

OpenStudy (alexandervonhumboldt2):

np